# Compression and Distillation of Data Quadruplets in Non-intrusive Reduced-order Modeling

### v3

@@ Abstract
This paper introduces a quadrature-free, data-driven approach to balanced truncation for both continuous-time and discrete-time systems. The method non-intrusively constructs reduced-order models using available transfer function samples from the right half of the $s$-plane. It is highlighted that the proposed data-driven balanced truncation and existing quadrature-based balanced truncation algorithms share a common feature: both compress their respective data quadruplets to derive reduced-order models. Additionally, it is demonstrated that by using different compression strategies, these quadruplets can be utilized to develop three data-driven formulations of the IRKA for both continuous-time and discrete-time systems. These formulations non-intrusively generate reduced models using transfer function samples from the $j\omega$-axis or the right half of the $s$-plane, or impulse response samples. Notably, these IRKA formulations eliminate the necessity of computing new transfer function samples as IRKA iteratively updates the interpolation points. The efficacy of the proposed algorithms is validated through numerical examples, which show that the proposed data-driven approaches perform comparably to their intrusive counterparts.
@@
ADI, Balanced truncation, Data-driven, $\mathcal{H}_{2}$-optimal, IRKA, , Non-intrusive
@@ 1 Introduction
The primary computational burden in BT lies in solving large-scale Lyapunov equations to compute the system Gramians. Various approaches, such as those mentioned in the surveys [5, 6], have been developed to efficiently compute these Gramians. These methods rely on the system’s explicit state-space representation, making BT an “intrusive” method. This is in contrast to “non-intrusive” methods, which are data-driven and depend solely on system response data—like transfer function samples or impulse response measurements—without requiring the system’s internal state-space representation [7, 8, 9, 10]. In [11], a non-intrusive BT algorithm based on numerical integration, called Quadrature-BT (QuadBT), is introduced. This algorithm constructs the ROM using transfer function samples at the $j\omega$-axis of the $s$-plane or samples of impulse response and its derivatives.
The $\mathcal{H}_{2}$-optimal MOR problem involves finding a local minimum for the (squared) $\mathcal{H}_{2}$ norm of the error transfer function. One of the key methods for achieving this local optimum is the Iterative Rational Krylov Algorithm (IRKA) [12]. A non-intrusive, data-driven version of IRKA was introduced in [13], based on the interpolatory framework proposed in [7]. This approach requires only transfer function samples and their derivatives to compute the local optimum, making it data-driven and non-intrusive. However, because IRKA is iterative, the sampling points are updated at each iteration and cannot be predetermined. Instead, IRKA identifies the optimal sampling points through successive iterations. If these new samples must be obtained experimentally, the algorithm must pause until the new data is available. This poses practical challenges, as it may be difficult or even impossible to conduct experiments to gather transfer function samples each time the sampling points are updated.
@@
Over the past two decades, the low-rank Alternating-direction Implicit (ADI) method has proven highly effective in reducing the computational cost of BT [16]. It is now one of the most widely used and efficient BT algorithms in the literature [17]. In this paper, we introduce a non-intrusive, data-driven implementation of the low-rank ADI-based BT that constructs the ROM from transfer function samples in the right-half of the $s$-plane. Unlike QuadBT, this approach does not rely on numerical integration. Additionally, we propose three non-intrusive, data-driven implementations for IRKA, tailored to the type of data available. In cases where transfer function samples along the $j\omega$ axis or impulse response measurements are accessible, we present numerical integration-based algorithms that do not require new transfer function samples as IRKA updates the sampling points. For scenarios where transfer function samples in the right-half of the $s$-plane are available, we propose a version that does not require numerical integration and new transfer function samples as IRKA updates the sampling points. Additionally, all these data-driven implementations (both BT and IRKA) for continuous-time systems are extended to discrete-time systems in this paper. It is also briefly highlighted that the implementations utilizing transfer function samples in the right half of the $s$-plane can also be executed using input-output data instead of relying solely on transfer function samples.
The remainder of the paper is structured as follows. Section 2 provides the necessary background on MOR and briefly reviews existing MOR algorithms most relevant to this work. The main contributions of this research begin in Section 3, where a data-driven, non-intrusive implementation of ADI-based low-rank BT is proposed. Section 4 presents three new data-driven implementations of IRKA, tailored to the type of available data. Section 5 introduces two quadrature-based data-driven implementations of IRKA for discrete-time systems. In Section 6, the PORK algorithm is extended to discrete-time systems. Building on this, Section 7 formulates a quadrature-free data-driven implementation of BT for discrete-time systems, while Section 8 develops a quadrature-free data-driven implementation of IRKA. Section 9 elaborates on the concepts of compression and distillation in the context of data-driven MOR. The performance of the proposed algorithms is evaluated in Section 10. Finally, the paper concludes in Section 11.
@@ 2.2 Iterative Rational Krylov Algorithm (IRKA) [12]
The necessary conditions for a local optimum of $||G(s)-G_{r}(s)||_{\mathcal{H}_{2}}^{2}$ are given by:
@@
Since the ROM $G_{r}(s)$ is initially unknown, IRKA uses fixed-point iterations starting from an arbitrary initial guess of the interpolation data to search for the local optimum. After each iteration, the interpolation data is updated as $\sigma_{i}=\mu_{i}=-\hat{\lambda}_{i}$, $b_{i}=\hat{r}_{i}$, and $c_{i}=\hat{l}_{i}^{*}$ until convergence is achieved. Upon convergence, a local optimum of $||G(s)-G_{r}(s)||_{\mathcal{H}_{2}}^{2}$ is achieved.
@@ 2.4 Interpolatory Loewner framework [7]
$\frac{c_{i}G(\sigma_{i})b_{j}-c_{i}G(\mu_{j})b_{j}}{\sigma_{i}-\mu_{j}}$ $\approx c_{i}G^{\prime}(\sigma_{j})b_{j},$
$\frac{\sigma_{j}c_{i}G(\sigma_{j})b_{j}-\mu_{i}c_{i}G(\mu_{i})b_{j}}{\sigma_{r}-\mu_{r}}$ $\approx c_{i}G(\sigma_{j})b_{j}+\sigma_{j}c_{i}G^{\prime}(\sigma_{j})b_{j}.$
@@ 2.5 Balanced Truncation (BT) [4]
[…]
@@
2.6 Data-driven Quadrature-based Balanced Truncation (QuadBT)[11]
Our presentation of QuadBT differs slightly from the original formulation in [11]. This choice of presentation aims to emphasize that QuadBT, like all the algorithms proposed in this paper, compresses and distills data quadruplets to construct the ROM. The concepts of compression and distillation in the context of data-driven MOR will be discussed in detail in Section 9.
@@ 3 Low-rank ADI-based Non-intrusive Balanced Truncation for Continuous-time Systems
In this section, we propose a non-intrusive, data-driven implementation of BT using transfer function samples from the right-half of the $s$-plane, as opposed to the $j\omega$-axis, which is used for QuadBT. We also briefly discuss how this implementation can be executed using time-domain input-output data without any modifications.
@@
[…]
The key distinction lies in where $X(s)$ is interpolated: in numerical integration, $X(s)$ is interpolated along the $j\omega$-axis, whereas in interpolatory projection methods like the ADI method, the interpolation occurs in the right-half of the $s$-plane.
In [20], it is shown that the low-rank approximation of Lyapunov equations produced by the block version of PORK is identical to that produced by the ADI method [16] when the mirror images of the interpolation points are used as ADI shifts. The block version of PORK enforces block interpolation instead of tangential interpolation. Over the past few decades, the ADI method has been highly successful in extending the applicability of BT to large-scale systems [17, 21]. In the sequel, a data-driven implementation of the block version of PORK-based BT is formulated, which produces results identical to the ADI-based BT.
@@
The pseudo-code for the data-driven ADI-based BT (DD-ADI-BT) is provided in Algorithm 1.
Algorithm 1 DD-ADI-BT
@@ 3.1 Avoiding the Computation of $Q_{s}$ and $P_{s}$
In practice, as the number of interpolation points increases, $Q_{s}$ and $P_{s}$ start losing numerical rank, making the computation of inverses $Q_{s}^{-1}$ and $P_{s}^{-1}$ highly ill-conditioned. This situation can be avoided if we make the choice of ADI shifts $-\sigma_{i}$ and $-\mu_{i}$ a bit restrictive. Let us represent $-\sigma_{i}$ and $-\mu_{i}$ as follows:
@@ 3.2 Extension to Tangential-ADI Method
3.3 DD-ADI-BT via Input-Output Time-domain data
The core computation in the DD-ADI-BT algorithm is step 1, which computes the Loewner quadruplet $(E_{s},A_{s},B_{s},C_{s})$. This can be computed from input-output time-domain data using the algorithm presented in [23] without any modifications. The other steps of the DD-ADI-BT algorithm remain the same. This extends the applicability of nonintrusive, data-driven BT to input-output time-domain data, complementing the other data types discussed earlier. For a detailed description of how the Loewner quadruplet $(E_{s},A_{s},B_{s},C_{s})$ is constructed from input-output time-domain data, see [23].
For the sake of completeness, we present a small illustrative example demonstrating how to implement DD-ADI-BT using input-output time-domain data. Although not mentioned in [23], the Loewner quadruplet $(E_{s},A_{s},B_{s},C_{s})$ can be constructed using a single signal generator instead of the two used in [23], as will be shown later. The following example uses the single signal generator-based approach proposed in [24] to construct the Loewner quadruplet $(E_{s},A_{s},B_{s},C_{s})$, thereby generalizing the algorithm in [24] for data-driven BT.
Illustrative Example: Consider the following second-order system:
$E$ $=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},$ $A$ $=\begin{bmatrix}-3.2556&1.1671\\ -1.1671&-0.4906\end{bmatrix},$
$B$ $=\begin{bmatrix}-1.1810\\ -0.1793\end{bmatrix},$ $C$ $=\begin{bmatrix}1.1810&-0.1793\end{bmatrix}.$
The ADI shifts for approximating the Gramians $P$ and $Q$ are selected as $(-2.6141,-1.1321)$. Since the same shifts are used for both Gramians, we require samples of $G(s)$ and its derivative $G^{\prime}(s)$ at $s=(2.6141,1.1321)$. The following signal generator is used to produce the input signal $u(t)$:
$\dot{w}(t)=Sw(t),\quad u(t)=Lw(t),$
where
$S=\begin{bmatrix}2.6141&5.2282&3.4406&3.4406\\ 0&2.6141&3.4406&3.4406\\ 0&0&1.1321&2.2642\\ 0&0&0&1.1321\end{bmatrix},\quad L=\begin{bmatrix}-2.2865\\ -2.2865\\ -1.5047\\ -1.5047\end{bmatrix}^{T},\quad w(0)=\begin{bmatrix}1\\ 2\\ 3\\ 4\end{bmatrix}.$
The eigenvalues of $S$ are $(2.6141,2.6141,1.1321,1.1321)$. The repeated eigenvalues arise because both $G(s)$ and $G^{\prime}(s)$ need to be sampled at $s=(2.6141,1.1321)$. The output of this signal generator is fed into the second-order system, and the resulting output is recorded. The MATLAB implementation of this experiment is provided in the appendix. From the system’s output response and the signal generator’s state trajectories, a fourth-order interpolant is estimated, which is then used to construct the Loewner quadruplet $(E_{s},A_{s},B_{s},C_{s})$ (see Appendix). Applying DD-ADI-BT to this quadruplet yields the following second-order model:
$E_{r}$ $=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},$ $A$ $=\begin{bmatrix}-3.2580&-1.1588\\ 1.1588&-0.4789\end{bmatrix},$
$B$ $=\begin{bmatrix}1.1813\\ -0.1777\end{bmatrix},$ $C$ $=\begin{bmatrix}-1.1813&-0.1777\end{bmatrix}.$
The Hankel singular values of this model are $(0.2141,0.0330)$, which are very close to the Hankel singular values $(0.2142,0.0328)$ of the original system.
4 Data-driven Implementations of IRKA for Continuous-time Systems
IRKA is highly effective for constructing $\mathcal{H}_{2}$-optimal ROMs through iterative refinement of interpolation data. However, its data-driven implementation poses a significant practical challenge. Each IRKA iteration updates the interpolation points, necessitating new measurements of $G(\sigma_{i})b_{i}$, $c_{i}G(\sigma_{i})$, and $c_{i}G^{\prime}(\sigma_{i})b_{i}$. As a consequence, the algorithm must be paused to conduct new experiments, making it unsuitable for practical applications. In this section, three non-intrusive, data-driven implementations of IRKA are proposed, which rely on existing available data instead of requiring new measurements of $G(\sigma_{i})b_{i}$, $c_{i}G(\sigma_{i})$, and $c_{i}G^{\prime}(\sigma_{i})b_{i}$ each time IRKA updates the interpolation data triplet $(\sigma_{i},b_{i},c_{i})$.
@@ 4.1 Using Available Frequency Response Data
In industries such as aerospace, defense, and automotive, frequency-domain data is collected to construct the Fourier transform $G(j\omega)$ by exciting systems at various frequencies $\omega$ rad/sec. This data plays a critical role in numerous analysis and design tasks, including system identification, control design, resonance frequency calculation, and vibration analysis, among others [25, 26, 27, 28, 29]. In this subsection, we demonstrate that this existing data is sufficient for non-intrusive data-driven implementation of IRKA.
@@
[…]
@@
It is now evident that IRKA can be implemented using frequency response data $G(j\omega_{i})$ and $G(j\nu_{i})$, eliminating the need for repeated measurements of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever IRKA updates $\sigma_{i}$. The pseudo-code for our proposed algorithm, called “frequency-domain quadrature-based IRKA (FD-Quad-IRKA)”, is provided in Algorithm 2.
@@
2:  while(relative change in $\lambda_{i}$ tol)
@@
7:  Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(-\lambda_{1},\cdots,-\lambda_{r})$; $[b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $[c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@
Range of Frequency Domain Sampling: Let us restrict the integral range of (38) from [,][$\infty$,$\infty$][ - ∞ , ∞ ] to $[-\nu,\nu]$ rad/sec. Then $V_{\nu}=V\Big{|}_{-\nu}^{\nu}$ solves the following Sylvester equation:
@@
Alternatively, the integration limits of the numerical quadrature rule can be mapped to [,][$\infty$,$\infty$][ - ∞ , ∞ ]. For instance, the integration limits $[-1,1]$ in the Gauss-Legendre quadrature rule can be mapped to [,][$\infty$,$\infty$][ - ∞ , ∞ ] using the following transformation:

### v4

@@ Abstract
This paper introduces a quadrature-free, non-intrusive approach to balanced truncation. The method non-intrusively constructs reduced-order models using available transfer function samples from the right half of the $s$-plane. It is highlighted that the proposed non-intrusive balanced truncation and existing quadrature-based balanced truncation algorithms share a common feature: both compress their respective data quadruplets to derive reduced-order models. Additionally, it is shown that by using different compression strategies, these quadruplets can be used to develop three non-intrusive formulations of the Iterative Rational Krylov Algorithm (IRKA). These formulations non-intrusively generate reduced models using transfer function samples from the $j\omega$-axis or the right half of the $s$-plane, or impulse response samples. Notably, these IRKA formulations eliminate the necessity of computing new transfer function samples as IRKA iteratively updates the interpolation points. The non-intrusive algorithms developed in this paper are also extended to discrete-time systems. The efficacy of the proposed algorithms is validated through numerical examples, which show that the proposed non-intrusive approaches perform comparably to their intrusive counterparts.
@@
ADI, Balanced truncation, Data-driven, $\mathcal{H}_{2}$-optimal, IRKA, Non-intrusive
@@ 1 Introduction
The primary computational burden in BT lies in solving large-scale Lyapunov equations to compute the system Gramians. Various approaches, such as those mentioned in the surveys [5, 6], have been developed to efficiently compute these Gramians. These methods rely on the system’s explicit state-space representation, making BT an “intrusive” method. This is in contrast to “non-intrusive” methods, which depend solely on system response data—like transfer function samples or impulse response measurements—without requiring the system’s internal state-space representation [7, 8, 9, 10]. In [11], a non-intrusive BT algorithm based on numerical integration, called Quadrature-BT (QuadBT), is introduced. This algorithm constructs the ROM using transfer function samples at the $j\omega$-axis of the $s$-plane or samples of impulse response and its derivatives.
The $\mathcal{H}_{2}$-optimal MOR problem involves finding a local minimum for the (squared) $\mathcal{H}_{2}$ norm of the error transfer function. One of the key methods for achieving this local optimum is the Iterative Rational Krylov Algorithm (IRKA) [12]. A non-intrusive version of IRKA was introduced in [13], based on the interpolatory framework proposed in [7]. This approach requires only transfer function samples and their derivatives to compute the local optimum, making it non-intrusive. However, because IRKA is iterative, the sampling points are updated at each iteration and cannot be predetermined. Instead, IRKA identifies the optimal sampling points through successive iterations. If these new samples must be estimated experimentally, the algorithm must pause until the new approximations of samples are available. This poses practical challenges, as it may be difficult or even impossible to conduct experiments to estimate transfer function samples each time the sampling points are updated.
@@
Over the past two decades, the low-rank Alternating-direction Implicit (ADI) method has proven highly effective in reducing the computational cost of BT [16]. It is now one of the most widely used and efficient BT algorithms in the literature [17]. In this paper, we introduce a non-intrusive implementation of the low-rank ADI-based BT that constructs the ROM from transfer function samples in the right-half of the $s$-plane. Unlike QuadBT, this approach does not rely on numerical integration. Additionally, we propose three non-intrusive implementations for IRKA, tailored to the type of data available. In cases where transfer function samples along the $j\omega$ axis or impulse response measurements are accessible, we present numerical integration-based algorithms that do not require new transfer function samples as IRKA updates the sampling points. For scenarios where transfer function samples in the right-half of the $s$-plane are available, we propose a version that does not require numerical integration and new transfer function samples as IRKA updates the sampling points. Additionally, all these non-intrusive implementations (both BT and IRKA) for continuous-time systems are extended to discrete-time systems in this paper.
The remainder of the paper is structured as follows. Section 2 provides the necessary background on MOR and briefly reviews existing MOR algorithms most relevant to this work. The main contributions of this research begin in Section 3, where a non-intrusive implementation of ADI-based low-rank BT is proposed. Section 4 presents three new non-intrusive implementations of IRKA, tailored to the type of available data. Section 5 introduces two quadrature-based non-intruisve implementations of IRKA for discrete-time systems. In Section 6, the PORK algorithm is extended to discrete-time systems. Building on this, Section 7 formulates a quadrature-free non-intrusive implementation of BT for discrete-time systems, while Section 8 develops a quadrature-free non-intrusive implementation of IRKA. Section 9 elaborates on the concepts of compression and distillation in the context of non-intrusive MOR. The performance of the proposed algorithms is evaluated in Section 10. Finally, the paper concludes in Section 11.
@@ 2.2 Iterative Rational Krylov Algorithm (IRKA) [12]
The necessary conditions for a local optimum of $G(s)-G_{r}(s)||_{\mathcal{H}_{2}}^{2}$ are given by:
@@
Since the ROM $G_{r}(s)$ is initially unknown, IRKA uses fixed-point iterations starting from an arbitrary initial guess of the interpolation data to search for the local optimum. After each iteration, the interpolation data is updated as $\sigma_{i}=\mu_{i}=-\hat{\lambda}_{i}$, $b_{i}=\hat{r}_{i}$, and $c_{i}=\hat{l}_{i}^{*}$ until convergence is achieved. Upon convergence, a local optimum of $G(s)-G_{r}(s)||_{\mathcal{H}_{2}}^{2}$ is achieved.
@@ 2.4 Interpolatory Loewner framework [7]
$\frac{c_{i}G(\sigma_{j})b_{j}-c_{i}G(\mu_{i})b_{j}}{\sigma_{j}-\mu_{i}}$ $\approx c_{i}G^{\prime}(\sigma_{j})b_{j},$
$\frac{\sigma_{j}c_{i}G(\sigma_{j})b_{j}-\mu_{i}c_{i}G(\mu_{i})b_{j}}{\sigma_{j}-\mu_{i}}$ $\approx c_{i}G(\sigma_{j})b_{j}+\sigma_{j}c_{i}G^{\prime}(\sigma_{j})b_{j}.$
@@ 2.5 Balanced Truncation (BT) [4]
$P$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(j\omega E-A)^{-1}BB^{T}(-j\omega E^{T}-A^{T})^{-1}\,d\omega,$ (14)
$Q$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(-j\omega E^{T}-A^{T})^{-1}C^{T}C(j\omega E-A)^{-1}\,d\omega.$ (15)
@@
2.6 Non-intrusive Quadrature-based Balanced Truncation (QuadBT)[11]
Our presentation of QuadBT differs slightly from the original formulation in [11]. This choice of presentation aims to emphasize that QuadBT, like all the algorithms proposed in this paper, compresses and distills data quadruplets to construct the ROM. The concepts of compression and distillation in the context of non-intrusive MOR will be discussed in detail in Section 9.
@@ 3 Low-rank ADI-based Non-intrusive Balanced Truncation for Continuous-time Systems
In this section, we propose a non-intrusive implementation of BT using transfer function samples from the right-half of the $s$-plane, as opposed to the $j\omega$-axis, which is used for QuadBT.
@@
$P\approx\frac{1}{2\pi}\int_{-\infty}^{\infty}\tilde{X}(j\omega)\,d\omega.$
The key distinction lies in where $X(s)$ is interpolated: in numerical integration, $X(s)$ is interpolated along the $j\omega$-axis, whereas in interpolatory projection methods like the ADI method, the interpolation occurs in the right-half of the $s$-plane. However, the essence of interpolation is the same in both methods—replacing $X(j\omega)$ with its approximation.
In [20], it is shown that the low-rank approximation of Lyapunov equations produced by the block version of PORK is identical to that produced by the ADI method [16] when the mirror images of the interpolation points are used as ADI shifts. The block version of PORK enforces block interpolation instead of tangential interpolation. Over the past few decades, the ADI method has been highly successful in extending the applicability of BT to large-scale systems [17, 21]. In the sequel, a non-intrusive implementation of the block version of PORK-based BT is formulated, which produces results identical to the ADI-based BT.
@@
The pseudo-code for the non-intrusive ADI-based BT (NI-ADI-BT) is provided in Algorithm 1.
Algorithm 1 NI-ADI-BT
@@ 3.1 Avoiding the Computation of $Q_{s}$ and $P_{s}$
In practice, as the number of interpolation points increases, $Q_{s}$ and $P_{s}$ start losing numerical rank, making the computation of inverses $Q_{s}^{-1}$ and $P_{s}^{-1}$ highly ill-conditioned. This situation can be avoided if we make the choice of ADI shifts $\sigma_{i}$ and $\mu_{i}$ a bit restrictive. Let us represent $\sigma_{i}$ and $\mu_{i}$ as follows:
@@ 3.2 Extension to Tangential-ADI Method
4 Non-intrusive Implementations of IRKA for Continuous-time Systems
IRKA is highly effective for constructing $\mathcal{H}_{2}$-optimal ROMs through iterative refinement of interpolation data. However, its non-intrusive implementation poses a significant practical challenge. Each IRKA iteration updates the interpolation points, necessitating new estimations of $G(\sigma_{i})b_{i}$, $c_{i}G(\sigma_{i})$, and $c_{i}G^{\prime}(\sigma_{i})b_{i}$. As a consequence, the algorithm must be paused to compute new estimations, making it unsuitable for practical applications. In this section, three non-intrusive implementations of IRKA are proposed, which rely on existing available data instead of requiring new estimations of $G(\sigma_{i})b_{i}$, $c_{i}G(\sigma_{i})$, and $c_{i}G^{\prime}(\sigma_{i})b_{i}$ each time IRKA updates the interpolation data triplet $(\sigma_{i},b_{i},c_{i})$.
@@ 4.1 Using Available Frequency Response Data
In industries such as aerospace, defense, and automotive, frequency-domain data is collected to construct the Fourier transform $G(j\omega)$ by exciting systems at various frequencies $\omega$ rad/sec. This data plays a critical role in numerous analysis and design tasks, including system identification, control design, resonance frequency calculation, and vibration analysis, among others [23, 24, 25, 26, 27]. In this subsection, we demonstrate that this existing data is sufficient for non-intrusive implementation of IRKA.
@@
$V$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(j\nu E-A)^{-1}BL_{b}(-j\nu I+S_{b})^{-1}d\nu,$ (38)
$W^{*}$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(-j\nu I+S_{c})^{-1}L_{c}C(j\nu E-A)^{-1}d\nu,$ (39)
@@
It is now evident that IRKA can be implemented using frequency response data $G(j\omega_{i})$ and $G(j\nu_{i})$, eliminating the need for repeated estimations of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever IRKA updates $\sigma_{i}$. The pseudo-code for our proposed algorithm, called “frequency-domain quadrature-based IRKA (FD-Quad-IRKA)”, is provided in Algorithm 2.
@@
2:while(relative change in $\lambda_{i}$ ¿ tol)
@@
7: Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(-\lambda_{1},\cdots,-\lambda_{r})$; $b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@
Range of Frequency Domain Sampling: Let us restrict the integral range of (38) from $\infty$$\infty$] to $\nu,\nu]$ rad/sec. Then $V_{\nu}=V\Big|_{-\nu}^{\nu}$ solves the following Sylvester equation:
@@
Alternatively, the integration limits of the numerical quadrature rule can be mapped to $\infty$$\infty$]. For instance, the integration limits [1,1][-1,1] in the Gauss-Legendre quadrature rule can be mapped to $\infty$$\infty$] using the following transformation:
@@
Avoiding the Samples of $G^{\prime}(s)$: To implement IRKA, transfer function samples and their derivatives at $r$ interpolation points are required. These can be computed by numerically integrating a single integral, though with some loss of accuracy compared to Algorithm 2, since single-sided moment matching is less accurate than double-sided moment matching. Let us define $S_{b}$ and $L_{b}$ as follows:
$S_{b}$ $=\text{blkdiag}\left(\begin{bmatrix}\sigma_{1}&1\\ 0&\sigma_{1}\end{bmatrix},\cdots,\begin{bmatrix}\sigma_{r}&1\\ 0&\sigma_{r}\end{bmatrix}\right)\otimes I_{m},$
$L_{b}$ $=\begin{bmatrix}1&0&\cdots&1&0\end{bmatrix}\otimes I_{m}.$ (45)
By solving the Sylvester equation (9), we obtain:
$V=\Big[(\sigma_{1}E-A)^{-1}B\quad-(\sigma_{1}E-A)^{-1}E(\sigma_{1}E-A)^{-1}B\quad\cdots$
$(\sigma_{r}E-A)^{-1}B\quad-(\sigma_{r}E-A)^{-1}E(\sigma_{r}E-A)^{-1}B\Big],$
as described in [15]. Consequently,
$CV=\begin{bmatrix}G(\sigma_{1})&G^{\prime}(\sigma_{1})&\cdots&G(\sigma_{r})&G^{\prime}(\sigma_{r})\end{bmatrix}$.
Thus, by setting $S_{b}$ and $L_{b}$ as in (45), we can compute:
$\begin{bmatrix}G(\sigma_{1})&G^{\prime}(\sigma_{1})&\cdots&G(\sigma_{r})&G^{\prime}(\sigma_{r})\end{bmatrix}\approx\begin{bmatrix}G(j\omega_{1})&\cdots&G(j\omega_{n_{p}})\end{bmatrix}\hat{V}_{r}.$